Circular No. 5686

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

SUPERNOVA 1989aa IN ANONYMOUS GALAXY C. Pollas, Observatoire de la Côte d’Azur, reports his discovery of a

probable supernova of mag ∼ 19.5 at α = 1^{h}16^{m}33^{s}.46, δ = +13^{o}00′25′′.9 (equinox 1950.0), 1′′.5 west and 4′′.3 north of the nucleus of an anonymous galaxy. The object was seen on technical pan films (limiting mag 21-22) on 1989 Dec. 28.8 and 31.7 UT. Nothing could be detected at the location on 1992 Dec. 13.8 and 14.8 UT.

SUPERNOVAE 1992br and 1992bs J. Maza, University of Chile; and M. Hamuy, Cerro Tololo Interamer-

ican Observatory, report the discovery of two supernovae on 20-min unfiltered IIa-O plates taken by G. Valladares with the CTIO Curtis Schmidt telescope. SN 1992br, of m_{pg} ∼ 18.5, was discovered by M. Wischnjewsky on an exposure on 1992 Dec. 27.09 UT and is 2′′.5 east and 7′′ south of the nucleus of an anonymous galaxy at α = 1^{h}43^{m}55^{s}.4, δ = −56^{o}20′57′′ (equinox 1950.0). SN 1992bs, of m_{pg} ∼ 18, was discovered by L. González on an exposure on Dec. 28.17 UT and is 9′′ west and 4′′ north of the nucleus of an anonymous spiral galaxy at α = 3^{h}27^{m}37^{s}.2, δ = −37^{o}26′38′′ (equinox 1950.0). The supernovae were confirmed by R. Aviles on CCD B and V images (obtained on Dec. 30.19 and 31.23, respectively) with the CTIO 0.9-m telescope. Spectrograms covering the wavelength range 320-750 nm, obtained on Dec. 31 with the CTIO 1.5-m telescope, show in each case the characteristic Si II (635.5 nm) feature of type Ia supernovae. The minimum of the Si II absorption yields an expansion velocity (with respect to the parent galaxy’s rest frame) of ∼ 10 000 km/s for SN 1992br and ∼ 9300 km/s for SN 1992bs. Spectra of the nuclei of the parent galaxies yield redshifts z = 0.088 and 0.064, respectively. Given the low signal-to-noise ratio of the spectrum it is difficult to estimate the age of SN 1992br accurately, but it is possible, from comparison with the type Ia supernova 1992A, to state that this object is not more than two weeks after maximum light. In the case of SN 1992bs, the spectrum shows absorptions of Si II (635.5, 564.0 and 545.4 nm) and only a moderate resemblance to the spectrum of the type Ia supernova 1990N taken seven days before maximum light; no match could be achieved between the spectrum of SN 1992bs and post-maximum spectra of other type Ia supernovae.

1993 January 1 Brian G. Marsden